a line passing through (3,15) and (4,y) ia parallel to a line with slope -13. What is the value of y?
Answer by vertciel(183)

Hello there,
Don't be intimidated by the variable! Simply substitute what you have as information into the slope formula:
Slope = (y2 - y1)/(x2 - x1)
-13 = (y - 15)/(4 - 3)
(y - 15)/1 = -13 <-- I rewrote the equation so it's easier to read.
y - 15 = -13
Think you can take it from here?